OpenStudy (anonymous):

The students in a high school increases by 6% pe r year. How long does i t take to double ? (plz provide work )

OpenStudy (anonymous):

How many students initially?

OpenStudy (anonymous):

doesnt say but it has something to do with exponential/log equations thingys..

OpenStudy (anonymous):

OK. Let's assume the initial number of students was x. 2x = x(1.06)^(t) You need to solve for 't'. Do you understand the equation?

OpenStudy (anonymous):

http://capybara.wikispaces.com/file/view/exp+log+model+worksheet+with+answers.pdf if u need the exact place it came from

OpenStudy (anonymous):

i just dont know how to do it lol even if i already hav the answer..

OpenStudy (anonymous):

OK so, 2x/x = (1.06)^(t) 2 = (1.06)^(t) log 2 / log 1.06 = t 11.89 years.

OpenStudy (anonymous):

thx much btw sorry for the confusion lol